# Communication-Efficient Federated Optimization over Semi-Decentralized Networks

### v2

@@ 1 Introduction
$\min_{\boldsymbol{x}\in\mathbb{R}^{d}}f(\boldsymbol{x})\coloneqq\frac{1}{n}\sum_{i=1}^{n}f_{i}(\boldsymbol{x}),\qquad\text{where}~{}f_{i}(\boldsymbol{x})\coloneqq\frac{1}{m}\sum_{\boldsymbol{z}\in\mathcal{D}_{i}}\ell(\boldsymbol{x};\boldsymbol{z}).$ (1)
@@
Given that the communication complexity depends on the trade-off between the communication rounds and the per-round cost, emerging works focus on heterogeneous communication over semi-decentralized networks, to gain the best from both agent-to-agent and agent-to-server communication [${}^{+}$21, WFC22, GSHG22]. Such semi-decentralized networks—consisting of a centralized server and a network of agents—widely exist in many applications, such as autonomous vehicles [CRW21], energy systems [NMTA21] and ML systems [${}^{+}$23]. It has been observed that heterogeneous communication largely alleviates the heavy network dependence of distributed learning and tackles the communication bottleneck of the server [${}^{+}$21]. However, to the best of our knowledge, all of them rely on the assumption of bounded data dissimilarity across agents and a complete characterization of the convergence behavior with respect to the network heterogeneity is still lacking. More detailed discussions are delegated to Section 1.2.
@@ 1.1 Our contributions
To fill the void, we propose a communication-efficient algorithm called PISCO, which incorporates gradient-tracking techniques [ZM10] and multiple local updates [${}^{+}$17] for solving federated nonconvex optimization over semi-decentralized networks modeled by a probabilistic connection model (shown in Figure 2). The highlights of our contributions are as follows.
@@
    We prove that PISCO converges at a rate of $O\left(1/\sqrt{nT_{o}K}\right)$ for sufficiently large $K$, where $K$ is the number of communication rounds and $T_{o}$ is the number of local updates. Our result does not impose the strong assumptions on data heterogeneity. Moreover, increasing the number of local updates accelerates the convergence over semi-decentralized networks.
@@
    We show that the communication heterogeneity offered by a semi-decentralized network largely alleviates the network dependency of communication overheads in decentralized networks via a few agent-to-server communication rounds. For large and sparse networks (i.e., the mixing rate $\lambda_{w}\to 0$), with a small probability $p=O\left(\sqrt{\lambda_{w}}\right)$ of agent-to-server communication, the network dependency improves from $O(\lambda_{w}^{-2})$ to $O(\lambda_{w}^{-1})$.
@@
    We conduct extensive numerical experiments on real-world datasets over various network topologies to show the superior communication efficiency of PISCO. The results substantiate the convergence speedup brought by multiple local updates and the robustness of PISCO to data heterogeneity and various topologies, even for locally disconnected networks.
Figure 2: The semi-decentralized communication protocol, where the server can be accessed with probability $p$ and agents can communicate with their neighbors whenever the server is not available. Here, dot lines represents the agent-to-server communication, while the solid ones are for agent-to-agent communication.
@@ 1.2 Related works
As the size of the networked system increases, there are considerable algorithms developed for solving distributed nonconvex optimization. Roughly speaking, they can be categorized into two classes — decentralized algorithms [NO09, LS16, HHZ17, ${}^{+}$18] where agents are only allowed to exchange information with neighbors, and federated algorithms [${}^{+}$20, ${}^{+}$20, WLC22] where agents are able to communicate with the server directly. Early attempts apply (stochastic) gradient descent to distributed optimization, which performs well in practice [NO09, ${}^{+}$20]. However, the dissimilarity among local objectives could degenerate the performance under heterogeneous data and thus requires additional assumptions like bounded gradient or diminishing step-sizes. To eliminate such strong assumptions, many following works [LS16, HHZ17, ${}^{+}$20] have been developed, including gradient tracking (GT) techniques. The key idea of GT is to utilize dynamic average consensus [ZM10] for global gradient estimation, which has been incorporated with many distributed optimization algorithms to achieve faster convergence rates in nonconvex settings [SS19, SLH20, SMSC22, ${}^{+}$]. Our proposed PISCO also takes advantage of GT to inherit its robustness against data heterogeneity.
@@
Communication efficiency is one of the most important bottlenecks in distributed ML. In the decentralized setting, the communication complexity largely depends on the network topology, i.e., poor connectivity slows down the information mixing and thus requires more communication rounds to consensus [NOR18]. As for the federated ML, the communication burden of the centralized server may be unaffordable. To overcome such bottlenecks, a number of strategies are proposed for improving communication efficiency [${}^{+}$23], including: 1) compression methods: compressing the information for communication [KLSJ20, ${}^{+}$]; 2) multiple local communication and updates: executing multiple gossip communication [LCCC20, ${}^{+}$21] or successive local updates within one communication round [NAYU22, GC23, ${}^{+}$20, ${}^{+}$20], and 3) heterogeneous communication model over semi-decentralized networks: allowing both agent-to-agent and agent-to-server communication [${}^{+}$21, ${}^{+}$23, ${}^{+}$21]. In this paper, we aim to gain the best communication efficiency from probabilistic heterogeneous communication over semi-decentralized networks and enable multiple local updates for more communication savings.
@@
As mentioned above, semi-decentralized ML, resorting to heterogeneous communication, tackles both the communication bottleneck of the centralized server in federated ML and the heavy network dependency of decentralized ML. We mainly summarize the works in nonconvex setting that are mostly related to this paper, while referring readers to [${}^{+}$21, ${}^{+}$23] for the (strongly) convex setting. For nonconvex problems, Gossip-PGA [${}^{+}$21] first integrates Gossip SGD [NO09] with periodical global averaging and obtains a better scalability. It shows that intermittently communicating with the server can largely alleviate the heavy dependence on the network connectivity, especially for large or sparse networks. However, the theoretical results depend on the assumption of bounded similarity between local objectives. Moreover, HL-SGD [GSHG22] extends Gossip-PGA to the hierarchical networked structure with multiple clusters, while HA-Fed [WFC22] can be viewed as HL-SGD with momentum. Both of them enable intra-cluster gossip averaging and inter-cluster averaging, but also rely on the data heterogeneity assumptions which may be impractical in many real-world applications. Noted that all of them consider deterministic heterogeneous communication, i.e., agents/cluster can only communicate with the server every $H$ communication rounds, but the synchronization largely depends on the availability of the server. To this end, we consider the probabilistic communication model, where agents only exchange information with the server at the probability $p$. Furthermore, to the best of our knowledge, none of these approaches enables multiple local updates within a single communication round, whereas ours benefits from the linear speedup provided by the local updates. More detailed comparison can be found in Table 1.
Algorithm Convergence rate 11footnotemark: 1 Bounded dissimilarity Accessibility of server Multiple local updates
Mini-batch Large-batch
Gossip-PGA [${}^{+}$21] $O\left(\frac{1}{\sqrt{nK}}\right)$ $O\left(\left(\frac{G}{K}\right)^{\frac{2}{3}}\right)$ every $H$ rounds
Table 1: Comparison of ours and other semi-decentralized algorithms in nonconvex optimization, regarding to the convergence rate, algorithm design and data heterogeneity assumptions. Here, $K$ is the number of communication rounds, $n$ is the number of agents, $T_{o}$ is the number of multiple local updates within a single gossip/global communication round, $G$ is the quantity in bounded dissimilarity between local objective and the global objective, $G_{1}$ is the quantity in bounded intra-cluster dissimilarity as [GSHG22, Assumption 4], $G_{2}$ is the quantity in bounded inter-cluster dissimilarity as [GSHG22, Assumption 5].
@@ 1.3 Notation
Throughout this paper, we use the lowercase and uppercase boldface letters to represent vectors and matrices, respectively. We use $\|\boldsymbol{A}\|_{\mathsf{F}}$ for the Frobenius norm of a matrix $\boldsymbol{A}$, $\|\boldsymbol{A}\|_{2}$ for the largest singular value of a matrix $\boldsymbol{A}$, $\|\boldsymbol{a}\|_{2}$ for the $l_{2}$ norm of a vector $\boldsymbol{a}$, and $\otimes$ for the Kronecker product. In addition, we use $\boldsymbol{I}_{n}$ for the identity matrix of dimension $n$, $\boldsymbol{1}_{n}$ for the all-one vector of dimension $n$ and $\boldsymbol{O}_{d\times n}$ for the all-zero matrix of dimension $(d\times n)$. For any two real functions $f(\cdot)$ and $g(\cdot)$ defined on $\mathbb{R}^{+}$, $f(x)=O(g(x))$ if there exists a positive real constant $M$ and $x_{0}$ such that $f(x)\leq Mg(x)$ for any $x\geq x_{0}$. Note that $\leq$” can be interpreted in an element-wise fashion, if it is applied to vectors or matrices with the same dimension.
@@ 2 Preliminaries
Communication graphs.
@@
Mixing.
For any communication graph $\mathcal{G}$, the mixing of local communication can be formally characterized by the mixing matrix $\boldsymbol{W}=[w_{ij}]_{1\leq i,j\leq n}$ defined in Definition 1.
@@
Given an undirected communication graph $\mathcal{G}$, a nonnegative matrix $\boldsymbol{W}\in\mathbb{R}^{n\times n}$ is the mixing matrix, whose element $w_{ij}=0$ if and only if $\{i,j\}\notin\mathcal{E}$ and $i\neq j$ and $\boldsymbol{W}$ is doubly stochastic, i.e.,
$\boldsymbol{W1}_{n}=\mathbf{1}_{n}~{}and~{}\mathbf{1}^{\top}_{n}\boldsymbol{W}=\mathbf{1}_{n}^{\top}.$
The mixing rate of $\boldsymbol{W}$ is a nonnegative constant, i.e.,
@@
Note that the doubly stochasticity implies $\|\boldsymbol{W}\|_{2}\leq 1$ and the mixing rate $\lambda_{w}=1-\lambda^{2}\in[0,1]$, where $\lambda$ denotes the second largest eigenvalue. The mixing rate can depict the connectivity of the communication graph $\mathcal{G}$, or to say, the speed of information mixing. Mathematically, for any vector $\boldsymbol{x}\in\mathbb{R}^{d}$,
$\|\boldsymbol{W}\boldsymbol{x}-\overline{\boldsymbol{x}}\|_{2}^{2}\leq(1-\lambda_{w})\|\boldsymbol{x}-\overline{\boldsymbol{x}}\|_{2}^{2},$
@@
Stochastic gradient methods.
To improve the computational efficiency, one popular approach is to apply stochastic gradient on the mini-batch data samples, instead of using the full-batch gradient. In distributed setting, we define the local stochastic gradient for each agent $i\in[n]$ as:
$\boldsymbol{g}_{i}=\frac{1}{b}\sum_{\boldsymbol{z}_{i}\in\mathcal{Z}_{i}}\nabla\ell(\boldsymbol{x_{i}};\boldsymbol{z}_{i}),$ (2)
where $\mathcal{Z}_{i}\subset\mathcal{D}_{i}$ denotes the sampled data batch with size $|\mathcal{Z}_{i}|=b<m$. In this paper, we assume that $\mathcal{Z}_{i}$ is drawn i.i.d. from $\mathcal{D}_{i}$ with the same mini-batch size $b$ for every agent $i\in[n]$ and the local stochastic gradient $\boldsymbol{g}_{i}$ is an unbiased estimate of $\nabla f_{i}(\boldsymbol{x}_{i})$, i.e.,
$\mathbb{E}[\boldsymbol{g}_{i}]=\nabla f_{i}(\boldsymbol{x}_{i}).$
Gradient-tracking techniques.
@@
To address this issue, gradient-tracking (GT) techniques [LS16, NOS17, QL18] have been proposed, which utilizes gossip communication for global gradient estimation leveraging dynamic average consensus [ZM10]. To improve computational efficiency, DSGT [PN21] incorporates GT with stochastic gradient methods. The updates at the $k$-th iteration are defined as: every agent $i\in[n]$ updates its optimization variable $\boldsymbol{x}_{i}^{k}$ and gradient-tracking variable $\boldsymbol{y}_{i}^{k}$ by
@@
where $\eta>0$ is the step size, the initialization $\boldsymbol{y}_{i}^{0}=\boldsymbol{g}_{i}^{0}$, and $\boldsymbol{g}_{i}^{k}=\frac{1}{b}\sum_{\boldsymbol{z}_{i}^{k}\in\mathcal{Z}_{i}^{k}}\nabla\ell(\boldsymbol{x}_{i}^{k};\boldsymbol{z}_{i}^{k})$ represents the local stochastic gradient over the sampled data batch $\mathcal{Z}_{i}^{k}$ drawn i.i.d. from $\mathcal{D}_{i}$ with mini-batch size $b$. More recent works [LLKS23, GC23] prove that DSGT with multiple local updates is able to converge under high data heterogeneity in nonconvex setting.
3 Proposed Algorithm
In this section, we introduce PISCO, which exploits communication heterogeneity from the probabilistic communication model and inherits the robustness to data heterogeneity from GT. Before the depiction of PISCO, we first introduce some compact notation for convenience. Let $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$ and $\boldsymbol{y}_{i}\in\mathbb{R}^{d}$ denote the local optimization variable and gradient-tracking variable at every agent $i\in[n]$, respectively. Accordingly, we denote the collection of all the optimization variables and gradient-tracking variables by the matrix form
$\boldsymbol{X}=\left[\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{n}\right]\in\mathbb{R}^{d\times n}\quad\text{and}\quad\boldsymbol{Y}=\left[\boldsymbol{y}_{1},\boldsymbol{y}_{2},\ldots,\boldsymbol{y}_{n}\right]\in\mathbb{R}^{d\times n}.$
We define the distributed objective as $F(\boldsymbol{X})=\sum_{i=1}^{n}f_{i}(\boldsymbol{x}_{i})$, where the distributed gradient is denoted by
$\nabla F(\boldsymbol{X})=\left[\nabla f_{1}(\boldsymbol{x}_{1}),\nabla f_{2}(\boldsymbol{x}_{2}),\ldots,\nabla f_{n}(\boldsymbol{x}_{n})\right]\in\mathbb{R}^{d\times n}.$
Similarly, we denote the gradient of empirical loss given the sampling dataset $\mathcal{Z}=\{\mathcal{Z}_{i}\}_{i=1}^{n}$ as
$\nabla\ell(\boldsymbol{X};\mathcal{Z})=\left[\sum_{\boldsymbol{z}_{1}\in\mathcal{Z}_{1}}\nabla\ell(\boldsymbol{x}_{1};\boldsymbol{z}_{1}),\sum_{\boldsymbol{z}_{2}\in\mathcal{Z}_{2}}\nabla\ell(\boldsymbol{x}_{2};\boldsymbol{z}_{2}),\ldots,\sum_{\boldsymbol{z}_{n}\in\mathcal{Z}_{n}}\nabla\ell(\boldsymbol{x}_{n};\boldsymbol{z}_{n})\right]\in\mathbb{R}^{d\times n},$
where $\{\mathcal{Z}_{i}\}_{i=1}^{n}$ is randomly drawn from $\{\mathcal{D}_{i}\}_{i=1}^{n}$ with the mini-batch size $b$. With the local stochastic gradient defined in (2), the distributed stochastic gradient can be represented by

### v4

@@ 1 Introduction
$\min_{\boldsymbol{x}\in\mathbb{R}^{d}}f(\boldsymbol{x})\coloneqq\frac{1}{n}\sum_{i=1}^{n}f_{i}(\boldsymbol{x}),\;\text{where}~{}f_{i}(\boldsymbol{x})\coloneqq\frac{1}{m}\sum_{\boldsymbol{z}\in\mathcal{D}_{i}}\ell(\boldsymbol{x};\boldsymbol{z}).$ (1)
@@
Given that the communication complexity depends on the trade-off between the communication rounds and the per-round cost, emerging works focus on heterogeneous communication over semi-decentralized networks, to gain the best from both agent-to-agent and agent-to-server communication [CYZ+21, WFC22, GSHG22]. Such semi-decentralized networks—consisting of a centralized server and a network of agents—widely exist in many applications, such as autonomous vehicles [CRW21], energy systems [NMTA21] and ML systems [MSY+23]. It has been observed that heterogeneous communication largely alleviates the heavy network dependence of distributed learning and tackles the communication bottleneck of the server [CYZ+21]. However, to the best of our knowledge, all of them rely on the assumption of bounded data dissimilarity across agents and a complete characterization of the convergence behavior with respect to the network heterogeneity is still lacking. More detailed discussions on communication-saving strategies and semi-decentralized approaches are provided in Section 1.2.
Figure 2: The semi-decentralized communication protocol, where the server can be accessed with probability $p$ and agents can communicate with their neighbors whenever the server is not available. Here, dotted lines represents the agent-to-server communication, while the solid ones are for agent-to-agent communication.
@@ 1.1 Our contributions
To fill the void, we propose a communication-efficient algorithm called PISCO, which incorporates gradient-tracking techniques [ZM10] and multiple local updates [MMR+17] for solving federated nonconvex optimization over semi-decentralized networks modeled by a probabilistic connection model (shown in Figure 2). Such a semi-decentralized communication model (with local updates) allows PISCO to be viewed as a special form of gradient-tracking-based algorithms with time-varying networks. However, existing convergence guarantees for nonconvex optimization [LS16, LW20, HY22], cannot fully characterize the benefits of agent-to-server communication and multiple local updates. Specifically, applying previous analyses would yield convergence results that depend on the spectral gap of the least connected network—i.e., the underlying gossip communication network—while failing to capture the value of agent-to-server communication. To quantify these benefits, our analysis is of independent interest and can be readily extended to time-varying networks. The highlights of our contributions are as follows.
@@
    We prove that PISCO converges at a rate of $O\left(1/\sqrt{nT_{o}K}\right)$ for sufficiently large $K$, where $K$ is the number of communication rounds and $T_{o}$ is the number of local updates. Our result does not impose the strong assumptions on data heterogeneity. Moreover, increasing the number of local updates accelerates the convergence over semi-decentralized networks. See Table 1 for a detailed comparison with prior art.
@@
    We show that the communication heterogeneity offered by a semi-decentralized network largely alleviates the network dependency of communication overheads in decentralized networks via a few agent-to-server communication rounds. For large and sparse networks (i.e., the mixing rate $\lambda_{w}\to 0$), with a small probability $p=\Theta\left(\sqrt{\lambda_{w}}\right)$ of agent-to-server communication, the network dependency improves from $O(\lambda_{w}^{-2})$ to $O(\lambda_{w}^{-1})$.
@@
    We corroborate the superior communication efficiency of PISCO through simulations on real-world datasets. The results substantiate the convergence speedup brought by multiple local updates and the robustness of PISCO to data heterogeneity and various topologies, even for locally disconnected networks.
Algorithm Convergence rate  [^1] Here we only present the leading term of the rate for simplicity. Bounded data heterogeneity assumption Accessibility of server Multiple local updates
Mini-batch Large-batch
Gossip-PGA [CYZ+21] $O\left(\frac{1}{\sqrt{nK}}\right)$ $O\left(\left(\frac{G}{K}\right)^{\frac{2}{3}}\right)$ every $H$ rounds
Table 1: Comparison of ours and other semi-decentralized algorithms in nonconvex optimization using the same batch size, regarding the convergence rate, algorithm design and data heterogeneity assumptions. Here, $K$ is the number of communication rounds, $n$ is the number of agents, $T_{o}$ is the number of multiple local updates within a single gossip/global communication round, $G$ is the quantity in bounded dissimilarity between local objective and the global objective, $G_{1}$ is the quantity in bounded intra-cluster dissimilarity as [GSHG22, Assumption 4], $G_{2}$ is the quantity in bounded inter-cluster dissimilarity as [GSHG22, Assumption 5].
@@ 1.2 Related works
As the size of the networked system increases, there are considerable algorithms developed for solving distributed nonconvex optimization. Roughly speaking, they can be categorized into two classes — decentralized algorithms [NO09, LS16, HHZ17, TLY+18] where agents are only allowed to exchange information with neighbors, and federated algorithms [LHY+20, KKM+20, WLC22] where agents are able to communicate with the server directly. Early attempts apply (stochastic) gradient descent to distributed optimization, which performs well in practice [NO09, LHY+20]. However, the dissimilarity among local objectives could degenerate the performance under heterogeneous data and thus requires additional assumptions like bounded gradient or diminishing step-sizes. To eliminate such strong assumptions, many following works [LS16, HHZ17, KKM+20] have been developed, including gradient tracking (GT) techniques. The key idea of GT is to utilize dynamic average consensus [ZM10] for global gradient estimation, which has been incorporated with many distributed optimization algorithms to achieve faster convergence rates in nonconvex settings [SS19, SLH20, MSSC24, ZLL+22]. Our proposed PISCO also takes advantage of GT to inherit its robustness against data heterogeneity.
@@
Communication efficiency is one of the most important bottlenecks in distributed ML. In the decentralized setting, the communication complexity largely depends on the network topology, i.e., poor connectivity slows down the information mixing and thus requires more communication rounds to consensus [NOR18]. As for the federated ML, the communication burden of the centralized server may be unaffordable. To overcome such bottlenecks, a number of strategies are proposed for improving communication efficiency [CBD+23], including: 1) compression methods: compressing the information for communication [KLSJ20, ZLL+22]; 2) multiple local communication and updates: executing multiple gossip communication [LCCC20, CYZ+21] or successive local updates within one communication round [NAYU23, GC23, LHY+20, KKM+20]; 3) heterogeneous communication model over semi-decentralized networks: allowing both agent-to-agent and agent-to-server communication [CYZ+21, SSM+23, LHA+21]; and 4) adaptive communication strategies: utilizing event-triggered communication mechanisms [GG20, DYG+18, HYZ+23] and tailored adaptive communication topologies [SLJ+22, DJY+23, THTG24, LLWY24] for saving unnecessary communications. In this paper, we aim to gain the best communication efficiency from probabilistic heterogeneous communication over semi-decentralized networks and enable multiple local updates for more communication savings.
@@
As mentioned above, semi-decentralized ML, resorting to heterogeneous communication, tackles both the communication bottleneck of the centralized server in federated ML and the heavy network dependency of decentralized ML. We mainly summarize the works in nonconvex setting that are mostly related to this paper, while referring readers to [CYZ+21, PHC+23] for the (strongly) convex setting. For nonconvex problems, Gossip-PGA [CYZ+21] first integrates Gossip SGD [NO09] with periodical global averaging and obtains a better scalability. It shows that intermittently communicating with the server can largely alleviate the heavy dependence on the network connectivity, especially for large or sparse networks. However, the theoretical results depend on the assumption of bounded similarity between local objectives. Moreover, HL-SGD [GSHG22] extends Gossip-PGA to the hierarchical networked structure with multiple clusters, while HA-Fed [WFC22] can be viewed as HL-SGD with momentum. Both of them enable intra-cluster gossip averaging and inter-cluster averaging, but also rely on the data heterogeneity assumptions which may be impractical in many real-world applications. Noted that all of them consider deterministic heterogeneous communication, i.e., agents/cluster can only communicate with the server every $H$ communication rounds, but the synchronization largely depends on the availability of the server. To this end, we consider the probabilistic communication model, where agents only exchange information with the server at the probability $p$. Furthermore, to the best of our knowledge, none of these approaches enable multiple local updates within a single communication round, whereas ours benefits from the linear speedup provided by the local updates. More detailed comparison can be found in Table 1. Note that this comparison is based on the same batch size, while [CYZ+21] and [GSHG22] may use fewer batch data per round.
@@ 1.3 Notation
Throughout this paper, we use the lowercase and uppercase boldface letters to represent vectors and matrices, respectively. We use $\|\boldsymbol{A}\|_{\mathsf{F}}$ for the Frobenius norm of a matrix $\boldsymbol{A}$, $\|\boldsymbol{A}\|_{2}$ for the largest singular value of a matrix $\boldsymbol{A}$, $\|\boldsymbol{a}\|_{2}$ for the $l_{2}$ norm of a vector $\boldsymbol{a}$, and $\otimes$ for the Kronecker product. In addition, we use $\boldsymbol{I}_{n}$ for the identity matrix of dimension $n$, $\boldsymbol{1}_{n}$ for the all-one vector of dimension $n$ and $\boldsymbol{O}_{d\times n}$ for the all-zero matrix of dimension $(d\times n)$. For any two real functions $f(\cdot)$ and $g(\cdot)$ defined on $\mathbb{R}^{+}$, $f(x)=O(g(x))$ if there exists a positive real constant $M$ and $x_{0}$ such that $f(x)\leq Mg(x)$ for any $x\geq x_{0}$. Similarly, $f(x)=\Theta(g(x))$ if there exist positive real constants $M_{1},M_{2}$ and $x_{0}$ such that $M_{1}g(x)\leq f(x)\leq M_{2}g(x)$ for any $x\geq x_{0}$. Note that $\leq$” can be interpreted in an element-wise fashion, if it is applied to vectors or matrices with the same dimension.
@@ 2 Preliminaries
2.1 Communication graph and mixing matrix
@@
Moreover, for any communication graph $\mathcal{G}$, the mixing of local communication can be formally characterized by the mixing matrix $\boldsymbol{W}=[w_{ij}]_{1\leq i,j\leq n}$ defined in Definition 1.
@@
Given an undirected communication graph $\mathcal{G}$, a nonnegative matrix $\boldsymbol{W}\in\mathbb{R}^{n\times n}$ is the mixing matrix, whose element $w_{ij}=0$ if and only if $\{i,j\}\notin\mathcal{E}$ and $i\neq j$ and $\boldsymbol{W}$ is doubly stochastic, i.e., $\boldsymbol{W1}_{n}=\mathbf{1}_{n}~{}and~{}\mathbf{1}^{\top}_{n}\boldsymbol{W}=\mathbf{1}_{n}^{\top}$. The mixing rate of $\boldsymbol{W}$ is a nonnegative constant, i.e.,
@@
Note that the doubly stochasticity implies $\|\boldsymbol{W}\|_{2}\leq 1$ and the mixing rate $\lambda_{w}=1-\lambda^{2}\in[0,1]$, where $\lambda$ denotes the second largest eigenvalue. The mixing rate can depict the connectivity of the communication graph $\mathcal{G}$, or to say, the speed of information mixing. Mathematically,
$\|\boldsymbol{W}\boldsymbol{x}-\overline{\boldsymbol{x}}\|_{2}^{2}\leq(1-\lambda_{w})\|\boldsymbol{x}-\overline{\boldsymbol{x}}\|_{2}^{2},\quad\forall\boldsymbol{x}\in\mathbb{R}^{d},$
@@
2.2 Stochastic gradient methods
To improve the computational efficiency, one popular approach is to replace the full-batch gradients with stochastic gradients on the mini-batch data samples. In distributed setting, we define the local stochastic gradient for each agent $i\in[n]$ as:
$\boldsymbol{g}_{i}=\frac{1}{b}\sum_{\boldsymbol{z}_{i}\in\mathcal{Z}_{i}}\nabla\ell(\boldsymbol{x_{i}};\boldsymbol{z}_{i}),\quad\forall\boldsymbol{x}_{i}\in\mathbb{R}^{d},$ (2)
where $\mathcal{Z}_{i}\subset\mathcal{D}_{i}$ denotes the sampled data batch. Here, we assume that $\mathcal{Z}_{i}$ is drawn i.i.d. from $\mathcal{D}_{i}$ with the same mini-batch size $b<m$ for every agent $i\in[n]$ for simplicity, while using an adaptive batch size could be of interest for better controlling the variance of stochastic gradients [BBN18]. Note that the local stochastic gradient $\boldsymbol{g}_{i}$ is an unbiased estimate of $\nabla f_{i}(\boldsymbol{x}_{i})$, i.e.,
$\mathbb{E}[\boldsymbol{g}_{i}]=\nabla f_{i}(\boldsymbol{x}_{i}),\quad\forall\boldsymbol{x}_{i}\in\mathbb{R}^{d}.$
2.3 Gradient-tracking techniques
@@
To address this issue, gradient-tracking (GT) techniques [LS16, NOS17, QL18] have been proposed, which utilizes gossip communication for global gradient estimation leveraging dynamic average consensus [ZM10]. Recently, DSGT [PN21] incorporates GT with stochastic gradient methods for computational efficiency. The updates at the $k$-th iteration are defined as: every agent $i\in[n]$ updates its optimization variable $\boldsymbol{x}_{i}^{k}$ and gradient-tracking variable $\boldsymbol{y}_{i}^{k}$ by
@@
where $\eta>0$ is the step-size and the initialization $\boldsymbol{y}_{i}^{0}=\boldsymbol{g}_{i}^{0}$. In addition, [XKK20] incorporates GT with variance-reduced techniques and [LCCC20] develops an approximate Newton-type methods with variance-reduced GT to further accelerate the convergence. More recent works [LLKS24, GC23] prove that DSGT with multiple local updates is able to converge under high data heterogeneity in nonconvex setting.
3 Proposed PISCO Algorithm
In this section, we introduce PISCO, which exploits communication heterogeneity from the probabilistic communication model and inherits the robustness to data heterogeneity from GT. Before the depiction of PISCO, we first introduce some compact-form notations for convenience. Let the matrices $\boldsymbol{X}=\left[\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{n}\right]\in\mathbb{R}^{d\times n}$ and $\boldsymbol{Y}=\left[\boldsymbol{y}_{1},\boldsymbol{y}_{2},\ldots,\boldsymbol{y}_{n}\right]\in\mathbb{R}^{d\times n}$ represent the collection of all the optimization variables and gradient-tracking variables, respectively. We also denote the gradient of empirical loss given the sampled batch dataset $\mathcal{Z}=\{\mathcal{Z}_{i}\}_{i=1}^{n}$ as
$\nabla\ell(\boldsymbol{X};\mathcal{Z})=\left[\sum_{\boldsymbol{z}_{1}\in\mathcal{Z}_{1}}\nabla\ell(\boldsymbol{x}_{1};\boldsymbol{z}_{1}),\ldots,\sum_{\boldsymbol{z}_{n}\in\mathcal{Z}_{n}}\nabla\ell(\boldsymbol{x}_{n};\boldsymbol{z}_{n})\right].$
With the local stochastic gradient as (2) in hand, the distributed stochastic gradient can be represented by